# EvRT-DETR: Latent Space Adaptation of Image Detectors for Event-based Vision

### v2

@@ 3.2 I2EvDet: Image-to-Event Detection Framework
We introduce I2EvDet (Image-to-Event Detection), a principled framework for adapting image-based object detectors to effectively process temporal data. While specialized architectures are typically developed for event-based object detection, our approach demonstrates that strategic adaptation of mainstream detectors can achieve superior performance with minimal architectural modifications.
Our I2EvDet framework applies latent space adaptation principles [16, 14, 17], which enable domain adaptation through minimal changes to pre-trained models. Our framework follows a two-stage process:
First, we establish a strong foundation by training the RT-DETR model on individual EBC frames (subsection 3.1), treating them as conventional images. This creates a detector with powerful spatial representation capabilities optimized for event data.
In the second stage, we transform this image-based detector into a video-capable model. We freeze the pre-trained RT-DETR parameters and insert lightweight RNN modules into the latent space of the encoder. The resulting model, which we call EvRT-DETR (Event RT-DETR), captures temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead.
This modular approach offers several methodological advantages. The architecture enforces an explicit disentanglement between spatial and temporal feature processing, facilitating a hierarchical optimization regime where each component operates within its disentangled parameter subspace. By constraining the RT-DETR parameters after initial convergence, we preserve the integrity of the learned latent manifold during temporal adaptation, mitigating potential representation collapse that could occur during end-to-end optimization. Moreover, this design maintains the architectural factorization inherent in RT-DETR, enabling computational flexibility through selective module deactivation when temporal reasoning is unnecessary for inference scenarios.
Figure 4: Temporal Adaptation Benefits of EvRT-DETR. Visualization of object detection across sequential 1Mpx frames captured by a vehicle-mounted camera that gradually comes to a stop. While the camera is in motion (left frames), both detectors maintain accurate bounding boxes of stationary vehicles. As camera motion decreases (middle frames), event data becomes increasingly sparse. When the camera stops completely (right frames), RT-DETR (top row) loses detection capability while EvRT-DETR (bottom row) maintains consistent detection by leveraging its temporal memory module. This demonstrates how our temporal adaptation effectively addresses a fundamental limitation of image object detection methods when applied to event-based cameras.
@@ 3.3 Temporal Modeling Considerations
A key theoretical consideration in our adaptation framework is the choice of temporal memory mechanism. Event cameras present a unique challenge: static objects become “invisible” as they generate no events when motionless. Their detection requires having a persistent memory of past events. While Transformers have dominated recent vision architectures, their fixed context windows limit temporal memory. We selected RNNs for their theoretically unbounded memory capacity through recurrent state, making them uniquely suited for maintaining object persistence during periods of invisibility in the event stream. This theoretical foundation informed our EvRT-DETR implementation..
@@ 3.4 Technical Implementation of EvRT-DETR
Our EvRT-DETR architecture implements the I2EvDet framework through strategic, minimal modifications to the RT-DETR model. As shown in Figure 3, we identify key integration points within the pre-trained model’s latent space where temporal information can be effectively incorporated without disrupting spatial feature extraction.
@@
Following latent space adaptation principles, we introduce specialized temporal processing modules at each scale, implementing three parallel RNN-based memory modules $\{\mathbf{R}_{3},\mathbf{R}_{4},\mathbf{R}_{5}\}$. This multi-scale temporal adaptation design ensures that temporal dynamics are captured at different levels of feature abstraction. Critically, each RNN module interfaces with its corresponding encoder feature map through residual connections, mathematically expressed as:
@@
~$\tilde{\mathcal{E}}_{i}^{t}$ $=\mathcal{E}_{i}^{t}+\alpha_{i}\cdot\mathcal{O}_{i}^{t},$ (4)
where $W_{i}^{down}\in\mathbb{R}^{C_{i}^{RNN}\times C_{i}}$ and $W_{i}^{up}\in\mathbb{R}^{C_{i}\times C_{i}^{RNN}}$ are the projection matrices that transform features into and out of the temporal processing space, $\mathcal{M}_{i}^{t}\in\mathbb{R}^{C_{i}^{RNN}\times H_{i}\times W_{i}}$ represents the temporal memory state for the $i$-th feature scale at time $t$, $\mathcal{O}_{i}^{t}$ denotes the corresponding temporal adaptation output, and $\alpha_{i}$ is a learnable scaling factor that modulates the influence of temporal information on the spatial representation, inspired by the ReZero technique [1]. In our implementation, the input projection matrix $W_{i}^{down}$ is integrated within the RNN architecture, which already handles dimensionality reduction through its internal gating mechanisms, while the output projection $W_{i}^{up}$ is explicitly implemented to ensure proper integration back into the frozen RT-DETR architecture. The transformed feature representation ~$\tilde{\mathcal{E}}_{i}^{t}$ thus incorporates temporal context while preserving the integrity of the original spatial representation $\mathcal{E}_{i}^{t}$, with our ablation studies showing that $C_{i}^{RNN}$ can be reduced 4x relative to $C_{i}$ with a minimal performance impact.
For the specific implementation of the RNN modules, we employ the ConvLSTM architecture [31], which effectively balances representational capacity with computational efficiency. This choice was informed by ConvLSTM’s demonstrated effectiveness in handling spatio-temporal data in the RVT model [11], allowing our adaptation to leverage established temporal modeling techniques.
@@ 4 Evaluation of the EvRT-DETR Model
Table 1: Performance Comparison on Event-Based Camera Datasets. Our approach achieves state-of-the-art results on both Gen1 and 1Mpx benchmarks, with significant improvements over previous methods. The first stage of our framework (RT-DETR-T/B) already achieves competitive performance using only image-based detection, while the proposed I2EvDet adapted models (EvRT-DETR-T/B) surpass all existing approaches. Runtime measurements are reported in milliseconds on an NVIDIA T4 GPU; values in parentheses represent original reported times on different hardware, with T4-equivalent times calculated using FLOPS/throughput conversion.
We evaluate our EvRT-DETR model on two standard EBC object detection benchmarks: Gen1 [6] and 1Mpx [26] (also known as Gen4) datasets. These datasets contain diverse driving scenarios captured by Prophesee’s neuromorphic vision sensors, providing an ideal testbed for validating our adaptation approach under real-world conditions.
@@ 4.1 Experimental Setting
We evaluate on two standard Prophesee automotive datasets: Gen1 [6] and 1Mpx [26]. Gen1 contains 39+ hours of recordings at $304\times 240$ resolution with manual annotations (1-4 Hz) for cars and pedestrians, presenting challenges with sparse labels and stationary objects. 1Mpx offers 14+ hours at higher resolution ($1280\times 720$) with denser automatic annotations (60 Hz) for cars, pedestrians, and two-wheelers. Both datasets are organized as 60-second clips with standard train/test splits. 1Mpx’s higher resolution and frame rate enables evaluation on more complex scenes, while Gen1’s annotation pattern tests robustness to sparse temporal supervision.
@@
[…]
For the Gen1 dataset (original size $(240,304)$), we pad frames to $(256,320)$ to ensure divisibility by 32, as required by the feature pyramid architecture of RT-DETR. For 1Mpx, we downsize from $(720,1280)$ to $(360,640)$ using bilinear interpolation and pad to $(384,640)$. This maintains consistency with prior work while allowing fair computational comparison.
@@
We train EvRT-DETR in two stages: (1) training the RT-DETR detector on individual event frames using standard practices (Adam optimizer, EMA of weights), then (2) freezing these parameters and training only the lightweight RNN modules on both random and sequential clips to develop temporal memory. The first stage establishes strong spatial feature extraction, while the second adds temporal processing with minimal parameter overhead. For temporal training, we maintain RNN memory continuity across sequential clips while resetting for random clips, using different clip lengths for Gen1 (21 frames) and 1Mpx (10 frames). Full implementation details are provided in the supplementary material.
Evaluation Protocol
We evaluate EvRT-DETR using the standard COCO (Common Objects in Context) Mean Average Precision (mAP) measure [20], which allows direct comparison with existing methods. For consistency with previous works, we use the EBC-specific implementation of the COCO metrics provided by Prophesee’s Automotive Dataset Toolbox [27, 26]. All inference times reported in Table 2 are standardized to NVIDIA T4 GPU performance for fair comparison.
@@ 4.2 Performance Analysis
Table 1 presents a comprehensive comparison of our EvRT-DETR models against state-of-the-art approaches on the Gen1 and 1Mpx benchmarks. The results can be analyzed in terms of different methodological approaches to event-based object detection:
First, our baseline adaptation using two RT-DETR variants—RT-DETR-T (Tiny, with ResNet-18 backbone) and RT-DETR-B (Base, with ResNet-50 backbone) - trained on simple event frames achieves competitive performance ($46.0\%$ and $47.6\%$ mAP on Gen1, $44.1\%$ and $45.2\%$ mAP on 1Mpx) without requiring specialized architectures or complex event representations. This demonstrates that modern mainstream object detectors can be effectively adapted to event data through appropriate training.
More importantly, our complete EvRT-DETR models with temporal adaptation significantly outperform all existing methods, achieving new state-of-the-art results on both benchmarks. EvRT-DETR-B reaches $52.7\%$ mAP on Gen1 ($+2.3\%$ over the previous best) and $50.1\%$ mAP on 1Mpx ($+1.4\%$ over the previous best). Even our lighter EvRT-DETR-T model surpasses all existing approaches while maintaining faster inference times.
As shown in Figure 4, the temporal adaptation module provides crucial capabilities for handling stationary objects - a fundamental challenge for event cameras. When objects are in motion, both RT-DETR and EvRT-DETR perform well, but when motion stops and event generation becomes minimal, EvRT-DETRs temporal memory enables consistent detection.
The performance improvements come with minimal computational overhead. While we focus primarily on detection accuracy in this work, the runtime measurements in Table 1 demonstrate that our approach maintains competitive inference speeds despite adding temporal processing capabilities.
@@ 4.3 Analysis of Adaptation Design Choices
We systematically analyze key design choices in our temporal adaptation module to better understand the factors contributing to EvRT-DETR’s performance. This analysis provides insights into efficient adaptation strategies for event-based object detection.
Model mAP (%) $\text{mAP}_{50}$ (%) $\text{mAP}_{75}$ (%)
EvRT-DETR-B (KS=1) $52.3$ $81.4$ $55.4$
EvRT-DETR-B (KS=3) $\mathbf{52.7}$ $\mathbf{82.0}$ $\mathbf{56.0}$
EvRT-DETR-B (KS=5) $52.0$ $81.3$ $54.8$
Table 2: Effect of ConvLSTM Kernel Size. Ablation study on Gen1 dataset showing optimal temporal adaptation performance with $3\times 3$ kernels
4.3.1 Spatial Context Size in Temporal Processing
While RVT [11] found optimal performance with $1\times 1$ ConvLSTM kernels (effectively pointwise LSTMs), our adaptation approach shows different optimal characteristics. As shown in Table 2, EvRT-DETR achieves best performance with $3\times 3$ kernels, suggesting that spatial context is valuable when adapting frozen RT-DETR features. This demonstrates how adaptation design choices may differ from specialized architectures built from the ground up. Increasing to $5\times 5$ kernels degrades performance, suggesting that the additional parameters and receptive field expansion do not provide beneficial information for the task given the available training data.
@@ 4.3.1 Spatial Context Size in Temporal Processing
Table 3: Effect of Temporal Module Placement. Performance on Gen1 dataset with ConvLSTM modules at different feature scales. The array notation $(x,y,z)$ indicates presence (1) or absence (0) of temporal modules at scales $\mathcal{E}_{3}$, $\mathcal{E}_{4}$, and $\mathcal{E}_{5}$ respectively. Low-level features benefit most from temporal adaptation.
4.3.2 Multi-scale Temporal Adaptation
A key aspect of our approach is the insertion of temporal modules at multiple scales in the feature hierarchy. Table 3 shows performance when selectively removing ConvLSTM cells from specific feature levels, with the array notation $(x,y,z)$ indicating presence (1) or absence (0) at each scale $\{\mathcal{E}_{3},\mathcal{E}_{4},\mathcal{E}_{5}\}$.
The results reveal a clear pattern: the lowest-level features ($\mathcal{E}_{3}$) benefit most significantly from temporal adaptation, with performance dropping substantially when this module is removed. The impact diminishes at higher feature levels, suggesting that temporal information is most valuable for fine-grained, lower-level features that capture detailed motion patterns. This insight can guide more efficient adaptation designs where computational resources are directed to where they provide maximum benefit.
@@ 4.3.2 Multi-scale Temporal Adaptation
Table 4: Impact of Temporal Module Capacity. Performance on Gen1 dataset with varying ConvLSTM hidden dimensions (M). The rightmost column shows the number of trainable parameters (in millions) in the temporal adaptation module. Minimal performance degradation at M=64 demonstrates efficient adaptation with reduced parameters, similar to LoRA-style approaches.

### v3

@@ 3.2 I2EvDet: Image-to-Event Detection Framework
I2EvDet is a principled framework for adapting image-based object detectors to process temporal data. While specialized architectures typically are developed for event-based object detection, this approach demonstrates that strategic adaptation of mainstream detectors can achieve superior performance with minimal architectural modifications.
The I2EvDet framework applies latent space adaptation principles [17, 15, 18], which enable domain adaptation through minimal changes to pre-trained models. The I2EvDet framework follows a two-stage process, and it is applicable to any object detection model with a clean separation between feature extraction and object detection modules (e.g., YOLO and DETR families).
First, we establish a solid foundation by training an object detector on individual EBC frames (subsection 3.1), treating them as conventional images. This creates a detector with powerful spatial representation capabilities optimized for event data.
In the second stage, we transform this image-based detector into a video-capable model. We freeze the pre-trained detector parameters and insert lightweight RNN modules between the feature extractor and detection components. We train the inserted temporal modules on the EBC video object detection task while keeping the original detector frozen. The resulting model captures temporal dynamics across frames while preserving the powerful representation capabilities of the original detector and adding minimal computational overhead.
This two-stage approach retains learned spatial representations of the original image detector while efficiently adding temporal capabilities. In this work, we apply I2EvDet to the RT-DETR model as the primary detector, resulting in the EvRT-DETR model. Additionally, we demonstrate the generalizability of the I2EvDet framework using experiments with YOLOX detectors.
Figure 4: Temporal Adaptation Benefits of EvRT-DETR. Visualization of object detection across sequential 1Mpx frames captured by a vehicle-mounted camera that gradually comes to a stop. While the camera is in motion (left frames), both detectors maintain accurate bounding boxes of stationary vehicles. As camera motion decreases (middle frames), event data becomes increasingly sparse. When the camera stops completely (right frames), RT-DETR (top row) loses detection capability, while EvRT-DETR (bottom row) maintains consistent detection by engaging its temporal memory module.
@@ 3.3 Temporal Modeling Considerations
A key theoretical consideration for this adaptation framework is the choice of temporal memory mechanism. Event cameras present a unique challenge: static objects become “invisible” as they generate no events when motionless. Their detection requires having a persistent memory of past events. While transformers have dominated recent vision architectures, their fixed context windows limit temporal memory. We employ RNNs for their theoretically unbounded memory capacity through recurrent state, making them uniquely suited for maintaining object persistence during periods of invisibility in the event stream. This theoretical foundation has informed the EvRT-DETR implementation.
@@ 3.4 Technical Implementation of EvRT-DETR
The EvRT-DETR architecture implements the I2EvDet framework via strategic, minimal modifications to the RT-DETR model. As shown in Figure 3, we identify key integration points within the pre-trained model’s latent space, where temporal information can be effectively incorporated without disrupting spatial feature extraction.
@@
Following latent space adaptation principles, we introduce specialized temporal processing modules at each scale, implementing three parallel RNN-based memory modules $\{\mathbf{R}_{3},\mathbf{R}_{4},\mathbf{R}_{5}\}$. This multiscale temporal adaptation design ensures that temporal dynamics are captured at different levels of feature abstraction. Critically, each RNN module interfaces with its corresponding encoder feature map through residual connections, mathematically expressed as:
@@
$\tilde{\mathcal{E}}_{i}^{t}$ $=\mathcal{E}_{i}^{t}+\alpha_{i}\cdot\mathcal{O}_{i}^{t},$ (4)
where $W_{i}^{down}\in\mathbb{R}^{C_{i}^{RNN}\times C_{i}}$ and $W_{i}^{up}\in\mathbb{R}^{C_{i}\times C_{i}^{RNN}}$ are the projection matrices that transform features into and out of the temporal processing space, $\mathcal{M}_{i}^{t}\in\mathbb{R}^{C_{i}^{RNN}\times H_{i}\times W_{i}}$ represents the temporal memory state for the $i$-th feature scale at time $t$, $\mathcal{O}_{i}^{t}$ denotes the corresponding temporal adaptation output, and $\alpha_{i}$ is a learnable scaling factor that modulates the influence of temporal information on the spatial representation inspired by the ReZero technique [1]. In our implementation, the input projection matrix $W_{i}^{down}$ is integrated within the RNN architecture, which already handles dimensionality reduction through its internal gating mechanisms. Meanwhile, the output projection $W_{i}^{up}$ is explicitly implemented to ensure proper integration back into the frozen RT-DETR architecture. Thus, the transformed feature representation $\tilde{\mathcal{E}}_{i}^{t}$ incorporates temporal context while maintaining the integrity of the original spatial representation $\mathcal{E}_{i}^{t}$. In addition, our ablation studies show that $C_{i}^{RNN}$ can be reduced 4x relative to $C_{i}$ with minimal performance impact.
For specific implementation of the RNN modules, we employ the ConvLSTM architecture [33], which effectively balances representational capacity with computational efficiency. This choice was informed by ConvLSTM’s demonstrated effectiveness in handling spatio-temporal data in the RVT model [12], allowing our adaptation to employ established temporal modeling techniques.
@@ 4 Evaluation of the EvRT-DETR Model
Table 1: Performance Comparison on Event-based Camera Datasets. Our approach achieves state-of-the-art results on both Gen1 and 1Mpx benchmarks with significant improvements over previous methods. The first stage of the framework (RT-DETR-T/B) already achieves competitive performance using only image-based detection, while the proposed I2EvDet-adapted models (EvRT-DETR-T/B) surpass all existing approaches. Runtime measurements are reported in milliseconds on an NVIDIA T4 GPU. Values in parentheses represent original reported times on different hardware with T4-equivalent times calculated using FLOPS/throughput conversion.
The EvRT-DETR model is evaluated on two standard EBC object detection benchmarks: Gen1 [6] and 1Mpx [28] (also known as Gen4) datasets. These datasets contain diverse driving scenarios captured by Prophesee’s neuromorphic vision sensors, providing an ideal testbed for validating our adaptation approach under real-world conditions.
@@ 4.1 Experimental Setting
We evaluate on two standard Prophesee automotive datasets: Gen1 [6] and 1Mpx [28]. Gen1 contains 39+ hours of recordings at $304\times 240$ resolution with manual annotations (1-4 Hz) for cars and pedestrians, presenting challenges with sparse labels and stationary objects. 1Mpx offers 14+ hours at higher resolution ($1280\times 720$) with denser automatic annotations (60 Hz) for cars, pedestrians, and two-wheelers. Both datasets are organized as 60-second clips with standard train/test splits. 1Mpx’s higher resolution and frame rate enable evaluation on more complex scenes, while Gen1’s annotation pattern tests robustness to sparse temporal supervision.
@@
We adopt the well-established stacked 2D Histogram approach from RVT protocols [12], constructing event frames from each $50\text{\,}\mathrm{m}\mathrm{s}$ time window. Each frame is subdivided into 10 equally time-spaced $5\text{\,}\mathrm{m}\mathrm{s}$ intervals, creating a representation with sufficient temporal resolution for our adaptation framework while remaining compatible with standard image processing architectures.
For the Gen1 dataset (original size $(240,304)$), we pad frames to $(256,320)$ to ensure divisibility by 32 as required by the feature pyramid architecture of RT-DETR. For 1Mpx, we downsize from $(720,1280)$ to $(360,640)$ using bilinear interpolation and pad to $(384,640)$. This maintains consistency with prior work while allowing for fair computational comparison.
@@
We train EvRT-DETR in two stages: (1) training the RT-DETR detector on individual event frames using standard practices (Adam optimizer; EMA of weights) then (2) freezing these parameters and training only the lightweight RNN modules on both random and sequential clips to develop temporal memory. The first stage establishes robust spatial feature extraction, while the second adds temporal processing with minimal parameter overhead. For temporal training, we maintain RNN memory continuity across sequential clips while resetting for random clips using different clip lengths for Gen1 (21 frames) and 1Mpx (10 frames). Full implementation details are provided in the supplementary material.
Evaluation Protocol.
We evaluate EvRT-DETR using the standard COCO (Common Objects in Context) Mean Average Precision (mAP) measure [21] that allows direct comparison with existing methods. For consistency with previous works, we use the EBC-specific implementation of the COCO metrics provided by Prophesee’s Automotive Dataset Toolbox [29, 28]. All inference times reported in Table 2 are standardized to NVIDIA T4 GPU performance for fair comparison. Inference times include model forward pass and post-processing but exclude event preprocessing, following standard reporting practices. When prior work reported only preprocessing-inclusive times, we use their reported values directly.
@@ 4.2 Performance Analysis
Table 1 presents a comprehensive comparison of our EvRT-DETR models against state-of-the-art approaches on the Gen1 and 1Mpx benchmarks.
First, our baseline adaptation using two RT-DETR variants, RT-DETR-T (Tiny, with ResNet-18 backbone) and RT-DETR-B (Base, with ResNet-50 backbone), and trained on simple event frames achieves competitive performance ($46.0\%$ and $47.6\%$ mAP on Gen1; $44.1\%$ and $45.2\%$ mAP on 1Mpx) without requiring specialized architectures or complex event representations. This demonstrates that modern mainstream object detectors can be effectively adapted to event data through appropriate training.
More importantly, our complete EvRT-DETR models with temporal adaptation significantly outperform all existing methods, achieving new state-of-the-art results on both benchmarks. EvRT-DETR-B reaches $52.7\%$ mAP on Gen1 ($2.3\%$ over the previous best) and $50.1\%$ mAP on 1Mpx ($1.4\%$ over the previous best). Even our lighter EvRT-DETR-T model surpasses all existing approaches while maintaining faster inference times.
Figure 4 shows how the temporal adaptation module provides crucial capabilities for handling stationary objects – a fundamental challenge for event cameras. When objects are in motion, both RT-DETR and EvRT-DETR perform well, yet, when motion stops and event generation becomes minimal, EvRT-DETRs temporal memory enables consistent detection.
These performance improvements come with minimal computational overhead. While we focus primarily on detection accuracy in this work, the runtime measurements in Table 1 demonstrate that our approach maintains competitive inference speeds despite adding temporal processing capabilities.
4.3 Framework Generalizability
Detector mAP (%)
Original I2EvDet Improvement
RT-DETR-T 46.0 52.3 +6.3+6.3
RT-DETR-B 47.6 52.7 +5.1+5.1
YOLOX-T 36.0 42.4 +6.4+6.4
YOLOX-X 43.4 47.8 +4.4+4.4
Table 2: Generalizability of I2EvDet Framework Across Different Object Detection Architectures on the Gen1 Dataset. Our temporal adaptation consistently improves performance regardless of the base detector architecture.
To demonstrate that our results are not specific to the RT-DETR architecture, we apply the I2EvDet framework to YOLOX [9]. YOLOX is a CNN-based detector that contrasts with RT-DETR’s transformer architecture and has been used in other EBC applications [42, 12]. Table 2 shows that our temporal adaptation approach consistently improves performance across different baseline architectures on the Gen1 dataset. The substantial improvements (4.4-6.4 mAP) across both RT-DETR and YOLOX variants validate that I2EvDet represents a general framework for adapting image-based detectors to temporal domains rather than an architecture-specific optimization. Additional details are provided in the supplementary material.
@@ 4.4 Analysis of Adaptation Design Choices
We systematically analyze key design choices in our temporal adaptation module to better understand the factors contributing to EvRT-DETR’s performance. This analysis provides insights into efficient adaptation strategies for event-based object detection. Additional ablation studies are provided in the supplementary material.
@@
Table 3: Effect of Temporal Module Placement. Performance on the Gen1 dataset with ConvLSTM modules at different feature scales. The array notation $(x,y,z)$ indicates presence (1) or absence (0) of temporal modules at scales $\mathcal{E}_{3}$, $\mathcal{E}_{4}$, and $\mathcal{E}_{5}$, respectively. Low-level features benefit most from temporal adaptation.
4.4.1 Multiscale Temporal Adaptation
A crucial aspect of our approach is the insertion of temporal modules at multiple scales in the feature hierarchy. Table 3 shows performance when selectively removing ConvLSTM cells from specific feature levels with the array notation $(x,y,z)$ indicating presence (1) or absence (0) at each scale $\{\mathcal{E}_{3},\mathcal{E}_{4},\mathcal{E}_{5}\}$.
The results reveal a clear pattern: the lowest-level features ($\mathcal{E}_{3}$) benefit most significantly from temporal adaptation as performance drops substantially when this module is removed. The impact diminishes at higher feature levels, suggesting that temporal information is most valuable for fine-grained, lower-level features that capture detailed motion patterns. This insight can guide more efficient adaptation designs, wherein computational resources are directed to where they provide maximum benefit.
@@ 4.4.1 Multiscale Temporal Adaptation
Table 4: Impact of Temporal Module Capacity. Performance on Gen1 dataset with varying ConvLSTM hidden dimensions (M). The rightmost column shows the number of trainable parameters (in millions) in the temporal adaptation module.